OpenStudy (anonymous):

Divide: 24u^7-21u^4 _______________ 3u^3

OpenStudy (anonymous):

I'll help gimme a sec

OpenStudy (anonymous):

I came up with \[8u^4-7u\]

OpenStudy (mertsj):

\[\frac{24u^7-21u^4}{3u^3}=\frac{24u^7}{3u^3}-\frac{21u^4}{3u^3}=8u^4-7u\]

OpenStudy (anonymous):

Yay I was right ^_^

OpenStudy (unklerhaukus):

\[\frac{24u^7-21u^4}{3u^3}=\frac{(3\times8\times u^{3+4})-(3\times7\times u^{3+1})}{3u^3}=\frac{3u^3(8u^4+7u)}{3u^3}\]

OpenStudy (anonymous):

I got it on my own, thank you.